OpenStudy (anonymous):

use the distance formula to find the distance between points (-4,-1) and (2,7).show all work please

OpenStudy (anonymous):

distance formula: sqrt [ (y_2 - y_1)^2 + (x_2 - x_1)^2) e.g. if we had points P(2,3) and Q(4,5), then sqrt [(3+5)^2 + (2+4)^2] = sqrt (64 + 36) = sqrt (100) = 10

OpenStudy (anonymous):

thanks